Energy provides an easy measure of the vital needs for us to survive over a wide variety of foods and nutrients. For historical reasons, a number of units have been developed and used to quantify energy, such as J, Cal, kCal, BTU. Suppose that a normal person consumes (and therefore emits) 2000 kCal of energy per day. We also know that in New England, it takes approximately 45 BTU of energy per day to keep one square foot of house warm. Without worrying about whether these people can fit in a house, how many people do you think it would require to keep an average sized house (2000 sqft) in New England warm solely using the energy these people emit? Show the steps of your calculations
Answer #1

45 BTU of energy per day is needed to keep one square foot house warm.

To keep 2000 square foot house warm; the required energy is:

(45 BTU/ 1 sqft)× 2000 sqft = 90000 BTU.

Now convert BTU to kcal. The conversion formula is:

1 Btu = 0.252 kcal.

So; 90000 Btu = (90000 × 0.252) kcal = 22680 kcal.

One person emits 2000 kcal of energy per day. So; no of persons required to emit 22680 kcal energy per day:

22680 kcal × 1 person/ 2000 kcal = 11.34 persons = 12 persons (approx)

So; 12 persons are required.